40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Oct 26 19:10:59 2015, gautam, Update, Green Locking, AUX PDH loop characterization 
    Reply  Tue Dec 8 23:24:08 2015, gautam, Update, Green Locking, Y end laser (Lightwave) PZT calibration Ycalib_8Dec.pdf
       Reply  Sun Dec 13 21:55:28 2015, gautam, Update, Green Locking, Y end laser (Lightwave) PZT calibration IMG_5972.JPGServoY_TF_13Dec2015.pdfDatanCode.zipPSD_916Hz.pdf
          Reply  Mon Dec 14 16:27:11 2015, gautam, Update, Green Locking, Y-end AUX PDH noise breakdown YAUX_NB_Dec2015.pdfPDH_errSig_Calib.pdf
             Reply  Mon Jan 4 16:45:11 2016, gautam, Update, Green Locking, Y-end AUX PDH noise breakdown ErrSigBreakdown.pdfcontrolSigBreakdown.pdfYEnd_PDH_OLTF.pdf
          Reply  Mon Jan 4 16:09:54 2016, gautam, Update, Green Locking, Y end laser (Lightwave) PZT calibration choosingExcFreqs.pdflaserPZTcalib.pdf
Message ID: 11879     Entry time: Mon Dec 14 16:27:11 2015     In reply to: 11877     Reply to this: 11907
Author: gautam 
Type: Update 
Category: Green Locking 
Subject: Y-end AUX PDH noise breakdown 

Summary:

I've attached the results from my measurements of the noise characteristics of the Y-end auxiliary PDH system.

Details:

The following spectra were measured, in the range DC-1MHz:

  1. Analyzer noise floor (measured with input terminated)
  2. Green REFL PD dark noise (measured with the Y-end green shutter closed)
  3. Mixer noise (measured with input to mixer terminated - measured with an SR560 with a gain of 100)
  4. Servo noise (measured with input to servo terminated)
  5. In loop error signal (measured with green locked to Y-arm, LSC off - using monitor point on PDH box)
  6. In loop control signal (measured with green locked to Y-arm, LSC off - using monitor point on PDH box)

In order to have good spectral resolution, the frequency range was divided into 5 subsections: DC-200Hz, 200Hz-3.4kHz, 3.4kHz-16.2kHz, 10kHz-100kHz, 100kHz-1MHz. The first three are measured using the SR785, while the last two ranges are measured with the Agilent network analyzer. The spectrum of the mixer output with its input terminated was quite close to the analyzer noise floor - hence, this was measured with an S560 preamplifier set to a gain of 100, and subsequently dividing the ASD by 100. To convert the Y-axis from V/rtHz to Hz/rtHz, I used two conversion factors: for the analyzer noise floor, PD dark noise, mixer noise and in-loop error signal, I made an Optickle simulation of a simple FP cavity (all parameters taken from the wiki optics page, except that I put in Yutaro's measured values for the arm loss and a modulation depth of 0.21 which I estimated as detailed here), and played around with the demodulation phase until I got an error signal that had the same qualitative shape as what I observed on an oscilloscope with the arms freely swinging (feedback to the laser PZT disabled). The number I finally used is 45.648 kHz/V (the main horns were 800mV peak-to-peak on an oscilloscope trace, results of the Optickle FP cavity simulation shown in Attachment #2 used to calibrate the X-axis). For the servo noise spectrum and in-loop control signal, I used the value of 2.43 MHz/V as determined here

I'm not sure what to make of the strong peaks in the mixer noise spectrum between ~60Hz and 10kHz - some of the more prominent peaks are 60Hz harmonics, but there are several peaks in between as well (these have been confusing me for some time now, they were present even when I made the measurement in this frequency range using the Agilent network analyzer. My plan is to repeat these measurements for the Xend now. 

Attachment 1: YAUX_NB_Dec2015.pdf  196 kB  | Hide | Hide all
YAUX_NB_Dec2015.pdf
Attachment 2: PDH_errSig_Calib.pdf  50 kB  | Hide | Hide all
PDH_errSig_Calib.pdf
ELOG V3.1.3-